Calorimeter Chemistry Question

A calorimeter contains 950.0 g water that changes temperature by 11.3 °C when 1.00 g of diesel is burned. What is the heat of combustion of 1.00 g of diesel, in kilojoules?

The specific heat of water is 4.18 J/(g·°C). Remember that combustion is an exothermic chemical process. Please make sure that the sign of your answer indicates this. Round your answer to the tenths place.
Answer #1

Answer: -44.8 kJ


-4.18 J/(g·°C) * 950.0g = 3,971 J/°C

-3,971 J/°C * 11.3 °C = 44,872.3 J

-44,872.3 J / 1,000 = -44.8 kJ

answered by: Guest